OpenStudy (anonymous):

a body is dropped from a tower hits an obstacle 20m below so that its speed is halved if the total time of flight is 4s the height of the tower above the ground is (g=10m/s^2)

OpenStudy (anonymous):

60

OpenStudy (ajprincess):

say the time taken to reach the obstacle is t1. 20=1/2*10*t1^2 4=t1^2 t1=2 here the initial velocity is zero. when it hits the obstacle say the velocity is v v=10*2 =20 since it is given that the velocity is halved after it hits the obstacle the new velocity is 10 it is given that the total time of flight is 4s so the remaining time is 2s say the remaining height to b traveld is h=10*2+1/2*10*4 =20+20 =40 total height=40+20 =60m

OpenStudy (anonymous):

thanzz